# PRISMM-Bench: A Benchmark of Peer-Review Grounded Multimodal Inconsistencies

### v2

@@ 4.3 User Study and Linguistic Bias Analysis
Setup. Eight non-author participants with at least PhD-level computer science research experience each answered ten randomly sampled Ident task questions: five in Focused context and five in Document context. For each question, participants first answered without context (question + answer options only), then with context. Focused context consisted of cropped images and/or text excerpts; Document context contains links to original PDFs. The survey was implemented via a custom web interface (cf. App. E).
@@
Probing Linguistic Bias. To confirm this effect generalizes beyond the user study subset, we re-evaluated the same four representative LMMs on the full Ident dataset under both Natural Language and JSON formats (Table 3). The same pattern holds: under natural language, models achieve inflated accuracies without context (e.g. 60.3% for Gemini 2.5 Pro) but performance drops toward chance under JSON. Correspondingly, $R$ increases under JSON for all models, showing that structured outputs suppress linguistic shortcuts and force models to rely more on visual evidence.
Table 3: Impact of answer representation on without-context performance and visual reliance. Accuracy is reported for the Ident task with Focused Context. $R$ is computed according to Eq. 1.
@@
InternVL3.5 8B 43.5 27.5 27.1 30.0
InternVL3.5 38B 50.4 33.1 24.4 41.1
Qwen 2.5 VL 72B 47.7 19.7 34.7 24.7
Gemini 2.5 Pro 60.3 51.1 36.6 51.9
@@ 5 Conclusions
Limitations. Our benchmark is limited in scope: it currently focuses on AI-domain papers from ICLR 2025 and emphasizes rejected submissions to capture unresolved errors. As a result, both the scale and domain coverage are restricted. Future work should expand to other fields and venues, and explore inconsistencies that may persist in accepted papers, offering a broader and more representative testbed.
@@ 6 Ethics Statement
This work introduces PRISMM-Bench, a benchmark for evaluating multimodal large language models (MLLMs) on scientific document understanding. In developing the benchmark, we exclusively use publicly available research papers from ICLR 2025, which are distributed under the Creative Commons Attribution 4.0 (CC-BY 4.0) license. This license explicitly permits redistribution, remixing, and adaptation of the material with proper attribution, and we ensure that all source materials are used in full compliance with these terms.
@@
In the appendix, we illustrate qualitative examples (App. A), including a dataset viewer and examples of text-table and figure-equation inconsistencies. We then provide a comprehensive list of assets (App. B), detailing the data sources, licenses, and models used, including both open-source and proprietary ones. The ablations section (App. C) explores the impact of rasterization resolution on model performance and analyzes the effectiveness of Chain-of-Thought (CoT) reasoning with a detailed case study. The dataset construction section (App. D.1) explains our refined methodology for sourcing, filtering, and annotating inconsistencies from scientific papers, including a discussion of the annotation criteria, a custom-built annotation tool, and dataset statistics. Next, we detail the process of generating LLM-based questions (App. D.4) for our benchmark tasks (Inconsistency Identification, Remedy, and Pair-Match), including our debiasing strategies using structured JSON representations. Finally, we provide details on the user study implementation (App. E), the full LLM prompts (App. F) used for various tasks, and screenshots (App. G and H) of our annotation and survey applications to provide a clear understanding of our methodology.
Appendix A Qualitative Examples and Dataset Viewer
Alongside the source code, the supplementary materials contain an annotation viewer, which can be used to visually explore out dataset in the web browser. The viewer can be launched by opening the index.html file inside the annotation_viewer folder (we recommend using Google Chrome).
We furthermore show qualitative examples of a text–table inconsistency (Fig. 3) and a figure-equation inconsistency (Fig. 4), together with their corresponding evaluation tasks.
@@ C.1 Impact of Rasterization Resolution
Low resolutions harm performance. Most models showed significant drops at 72 DPI, up to -20.1% for InternVL3.5 38B. Open-weight models were generally more vulnerable, though even Gemini 2.5 Pro declined by -3.3%. Suprisingly, VILA H, despite being trained for high-resolution inputs, showed a slight accuracy gain at this lower setting.
@@
Resolution sensitivity is architecture-specific. Overall, the assumption that higher resolution improves inconsistency detection does not hold universally. Performance varies with model design and pretraining data, and high-resolution training does not guarantee an edge in handling scientific inconsistencies. Careful DPI control is crtical for fair evaluation. In our benchmark, 144 DPI provides a practical balance of visual clarity, computational cost, and cross-model comparability.
C.2 Chain-of-Thought Reasoning
Reasoning variants perform better than non-reasoning counterparts and reach results comparable to much larger models. For instance, InternVL3.5 8B achieves an average of 37.6%, rivaling large open-weight models and surpassing several 72B non-reasoning models. We therefore study how reasoning-enabled models leverage chain-of-thought (CoT) to improve performance on PRISMM-Bench. To do so, we re-evaluated a selection of reasoning models on the Ident task with Focused Context with reasoning turned off and compared the performance.
Disabling reasoning reveals the critical role of CoT in detecting subtle inconsistencies. For example, GLM 4.5V drops from 55.7% to 46.6% (-16.3%), InternVL3.5 8B from 48.9% to 40.0% (-18.2%), and InternVL3.5 38B suffers the largest decline, from 57.4% to 37.8% (-34.0%).
Why Reasoning Helps. To understand these gains, we focused on inconsistencies where InternVL3.5 38B succeeded with reasoning but failed without. We identified three consistent patterns: (1) Structured input handling. Reasoning-enabled models interpret the JSON-formatted options in natural language, clarifying subtle distinctions without exploring linguistic biases (cf. low without-context performance for reasoning models in Sec. 4.3). (2) Cross-modal grounding. CoT traces show models explicitly reasoning over both text and visuals, breaking complex information into smaller units and reusing them later in the reasoning chain. (3) Concept linking. Reasoning enables models to connect fine-grained context with domain knowledge and abstract concepts, allowing stronger logical inference beyond surface pattern recognition.
Case Study. Fig. C.2 illustrates this effect with the Unique Successful Jailbreaks metric, which must be non-negative. The figure, however, contains error bars extending below zeros, resulting in an inconsistency.
@@ D.1 Review Sourcing
    Conference selection: We shifted to the more recent ICLR 2025, focusing on papers without author rebuttals. This ensured that flagged inconsistencies were more likely to remain in the available PDFs.
@@
This refinement reduced our initial pool of 75,550 ICLR 2025 reviews to 12,366 reviews. The LLM outputs were stored in structured JSON format, with each paper ID associated with a list of flagged inconsistencies.
@@ D.2 Annotation Process
The annotation resulted in 262 inconsistencies from 242 ICLR papers. The average page count of each PDF was 18 pages. A total of 19 papers (7.9%) had more than one inconsistency. The paper subjects were equally distributed across the range of topics for ICLR 2025444https://iclr.cc/Conferences/2025/CallForPapers, with (1) representation learning (14.5%), (2) transfer learning (11.2%), (3) datasets and benchmarks (10.8%) and generative models (9.5%) being the most frequent topics.
We identified 13 categories of inconsistencies based on the elements involved, with the distribution shown in Fig. 9. The most common cases were figure–text mismatches and intra-figure inconsistencies.
Figure 9: Distribution of inconsistency types. We identified 13 categories of inconsistencies based on the elements involved. The most common cases are figure-text mismatches and intra-figure (figure-only) inconsistencies

### v3

@@ 4.4 User Study and Linguistic Bias Analysis
Setup. Eight non-author participants with at least PhD-level computer science research experience each answered ten randomly sampled Ident task questions: five in Focused context and five in Document context. We provide a detailed analysis of the representativeness of this subset in App. E.2. For each question, participants first answered without context (question + answer options only), then with context. Focused context consisted of cropped images and/or text excerpts; Document context contains links to original PDFs. The survey was implemented via a custom web interface (cf. App. E.1).
@@
Probing Linguistic Bias. To confirm this effect generalizes beyond the user study subset, we re-evaluated the same four representative LMMs on the full Ident dataset under both Natural Language and JSON formats (Table 3). The same pattern holds: under natural language, models achieve inflated accuracies without context (e.g. 61.2% for Gemini 2.5 Pro) but performance drops toward chance under JSON. Correspondingly, $R$ increases under JSON for all models, showing that structured outputs suppress linguistic shortcuts and force models to rely more on visual evidence.
Table 3: Impact of answer representation on without-context performance and visual reliance. Accuracy is reported for the Ident task. $R$ is computed according to Eq. 1. using Focused context as the with-context baseline.
@@
InternVL3.5 8B 45.6 17.1 28.6 29.3
InternVL3.5 38B 52.9 22.5 26.3 38.1
Qwen 2.5 VL 72B 49.7 16.1 36.5 20.8
Gemini 2.5 Pro 61.2 43.8 37.8 45.2
@@ 5 Conclusions
Limitations. Our benchmark is limited in scope: it currently focuses on AI-domain papers from ICLR 2024 & 2025 and emphasizes rejected submissions to capture unresolved errors. As a result, both the scale and domain coverage are restricted. Future work should expand to other fields and venues, and explore inconsistencies that may persist in accepted papers, offering a broader and more representative testbed.
@@ 6 Ethics Statement
This work introduces PRISMM-Bench, a benchmark for evaluating multimodal large language models (MLLMs) on scientific document understanding. In developing the benchmark, we exclusively use publicly available research papers from ICLR 2024 & 2025, which are distributed under the Creative Commons Attribution 4.0 (CC-BY 4.0) license. This license explicitly permits redistribution, remixing, and adaptation of the material with proper attribution, and we ensure that all source materials are used in full compliance with these terms.
@@
In the appendix, we first discuss illustrative qualitative examples (App. A), including examples of text-table and figure-equation inconsistencies. We then provide a comprehensive list of assets (App. B), detailing the data sources, licenses, and models used, including both open-source and proprietary ones. The ablations section (App. C) explores the impact of rasterization resolution on model performance, extends the analysis of the effectiveness of Chain-of-Thought (CoT) reasoning with a detailed case study, and validates our choice of the multiple-choice question format against open-ended evaluation. The dataset construction section (App. D.1) explains our refined methodology for sourcing, filtering, and annotating inconsistencies from scientific papers, including a discussion of the annotation criteria, a custom-built annotation tool, and dataset statistics. Next, we detail the process of generating LLM-based questions (App. D.4) for our benchmark tasks (Inconsistency Identification, Remedy, and Pair-Match), including our debiasing strategies using structured JSON representations. Finally, we provide details on the user study implementation and representativeness (App. E.1), the full LLM prompts (App. F) used for various tasks, and screenshots (App. G and H) of our annotation and survey applications to provide a clear understanding of our methodology.
Appendix A Qualitative Examples
We show qualitative examples of a text–table inconsistency in Fig. 3 and a figure-equation inconsistency in Fig. 4, together with their corresponding evaluation tasks.
@@ C.1 Impact of Rasterization Resolution
Low resolutions harm performance. Most models showed significant drops at 72 DPI, up to -20.1% for InternVL3.5 38B. Open-weight models were generally more vulnerable, though even Gemini 2.5 Pro declined by -3.3%. Surprisingly, VILA HD, despite being trained for high-resolution inputs, showed a slight accuracy gain at this lower setting.
@@
Resolution sensitivity is architecture-specific. Overall, the assumption that higher resolution improves inconsistency detection does not hold universally. Performance varies with model design and pretraining data, and high-resolution training does not guarantee an edge in handling scientific inconsistencies. Careful DPI control is critical for fair evaluation. In our benchmark, 144 DPI provides a practical balance of visual clarity, computational cost, and cross-model comparability.
C.2 Case Study: Chain-of-Thought Reasoning
To illustrate the effect of chain-of-thought reasoning discussed in Section 4.3, we present a representative example in Fig. C.2. The figure reports results for Unique Successful Jailbreaks, a count-based metric that is strictly non-negative. However, some error bars extend below the zero line on the y-axis—an inconsistency that invalidates the figure.
@@ C.2 Case Study: Chain-of-Thought Reasoning
C.3 Validity of Multiple-Choice vs. Open-Ended Evaluation
While open-ended generation is closer to real-world usage, the choice of a multiple-choice question (MCQ) format for our benchmark was motivated by the need for a reliable, reproducible, and bias-controlled evaluation protocol. To validate this choice, we conducted additional experiments comparing our MCQ format with an open-ended evaluation using an LLM-as-a-judge approach.
Setup. We tested a representative subset of the identification task using open-ended responses from eight models. Their outputs were rated by three different judges—Gemini 2.5 Pro, GPT-5 (high), and GLM 4.5V 106B—on a Likert-5 scale measuring semantic alignment with the ground-truth inconsistency. Each judge was evaluated both with and without Focused Context.
Table 5: LLM-as-a-judge evaluation. Scores are average Likert ratings (1-5); percentages indicate scores >> 3. R denotes the model’s rank per judge.
Candidate Model Gemini 2.5 Pro GPT-5 GLM 4.5V 106B
w/o context context w/o context context w/o context context
Score R Score R Score R Score R Score R Score R
[…]
Analysis of Open-Ended Evaluation. As shown in Table 5, our open-ended evaluation revealed several critical limitations that undermine its suitability for a standardized benchmark. First, we observed significant inter-judge inconsistency, where LLM judges produced diverging scores and rankings for identical model outputs. For instance, Gemini 2.5 Pro and GPT-5 frequently disagreed on absolute scores, leading to rank shifts of two to three positions for certain models depending on the judge employed. Furthermore, GLM 4.5V 106B proved particularly unstable in its role as a judge, further complicating reproducibility.
A second challenge relates to non-deterministic scoring in proprietary reasoning models. Since a temperature of zero is not always available for models such as GPT-5, repeated evaluations of the same answer do not yield deterministic results. This introduces noise into the benchmarking process, which is especially problematic when performance differences between high-performing models are small. Finally, we identified a clear judge-model similarity bias, where judges tended to rate outputs from their own model families more favorably. This phenomenon, documented in prior evaluations such as MT Bench (Zheng et al., 2023) and AlpacaEval (Dubois et al., 2025), makes it difficult to disentangle true model capability from stylistic alignment with the judge.
Conclusion. In contrast to open-ended evaluation, the MCQ format remains more reproducible, providing deterministic scoring, controlled difficulty through distractors, and identical evaluation conditions without dependence on third-party proprietary judges. Combined with our JSON-based debiasing strategies, the MCQ format provides a robust and reliable framework for PRISMM-Bench.
@@ D.1 Review Sourcing
    Conference selection: We shifted our focus on papers without author rebuttals. This ensured that flagged inconsistencies were more likely to remain in the available PDFs.
@@
This refinement reduced our initial pool of 120,329 ICLR reviews to 18,009 reviews. The LLM outputs were stored in structured JSON format, with each paper ID associated with a list of flagged inconsistencies.
@@ D.2 Annotation Process
The annotation resulted in 384 inconsistencies from 353 ICLR papers. The average page count of each PDF was 16 pages. A total of 29 papers (7.6%) had more than one inconsistency. The paper subjects were equally distributed across the range of topics for ICLR [^4] https://iclr.cc/Conferences/2025/CallForPapers, with (1) representation learning (26.0%), (2) transfer learning (10.4%), (3) generative models (9.6%) and datasets and benchmarks (7.8%) being the most frequent topics.
We identified 15 categories of inconsistencies based on the elements involved, with the distribution shown in Fig. 9. The most common cases were figure–text mismatches and intra-figure inconsistencies.
Figure 9: Distribution of inconsistency types. We identified 15 categories of inconsistencies based on the elements involved. The most common cases are figure-text mismatches and intra-figure (figure-only) inconsistencies.
@@ E.1 Implementation of the User Study
E.2 Representativeness of the User Study Subset
To ensure that the 40 samples used in the user study are representative of the full benchmark, we clarify that they were randomly drawn from the full dataset without any filtering or cherry-picking. As for some models, such as InternVL3.5 38B and Qwen2.5 VL 72B, accuracies on this subset were higher than their full-benchmark performance, we assessed whether this deviation was expected by running a 100-trial simulation. In each trial, we repeatedly sampled 40 random instances and computed the resulting scores.
As shown in Table 6, the simulation results indicate that the user-study subset happened to be in the upper quartile of model difficulty for these two models—a comparatively “easier” slice of the benchmark for them. Such variation is expected with small sample sizes and does not reflect systematic bias in the subset selection. Importantly, this strengthens our human-LMM comparison: even on a subset where LMMs performed better than their typical accuracy, humans still outperformed the models by a clear margin under both focused and whole-document conditions. If the subset had been closer to the mean model difficulty, the human-model gap would have been even larger. Thus, our user study results likely understate the core conclusion that humans substantially outperform current LMMs at detecting multimodal scientific inconsistencies.
Table 6: Model Performance Comparison: User Study vs. Simulation
Model (Context) User Study Subset (Reported) Simulation Mean Simulation Q3 (75th Percentile) Human Perf. (Reference)
InternVL 3.5 38B (Focused) 71.1% 58.6% 65.0% 77.5%
InternVL 3.5 38B (Whole Document) 40.5% 30.6% 35.0% 65.0%
Qwen 2.5 VL 72B (Focused) 65.8% 50.5% 55.0% 77.5%
Qwen 2.5 VL 72B (Whole Document) 48.6% 35.9% 40.0% 65.0%